onsider the following table, which gives a security analysts expected return on two stocks for two particuliar market returns ket RefurnAggressive StockDefensive Stock 2.1% 6% 16 ? 6% 10 25 a. What are the betas ot the two stocks? (Round your answers to 2 decimal places.) Beta A Beta D b. what is the expected rate of return on each stock if the market return is equally ikely to be 6% or 16%? Round your answers to 2 decinal places Rate of return on A Rate of return on D c. If the T bill rate is 8%, and the market return is equally likely to be 6% or 16%, what are heal nasof he wo stoc s? Negative values should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to 2 decimal places.) Aipha A Alpha D

please help. thank you.
Answer #1

a) Beta is the measure of movement of a stock in relation to the market. So, beta can be calculated as follows in our case -

Betastock = Change in return of stock / Change in market return

BetaA = (25% - 2.1%) / (16% - 6%) = 2.29

BetaD = (10% - 3.6%) / (16% - 6 %) = 0.64

b) Probability of each market return = 1 / 2 or 0.5

Expected returnA = 0.5 x 2.1% + 0.5 x 25% = 13.55%

Expected returnB = 0.5 x 3.6% + 0.5 x 10% = 6.80%

c) Alphastock = Expected returnstock - Risk free rate - Betastock x (Expected returnmarket - Risk free rate)

Expected returnmarket = 0.5 x 6% + 0.5 x 16% = 11%

AlphaA = 13.55% - 8% - 2.29 x (11% - 8%) = (-)1.32%

AlphaA = 6.80% - 8% - 0.64 x (11% - 8%) = (-)3.12%